OpenStudy (amberlykhan):

If you have six peoplein a room, and each person shakes hands with every ither person exactly once, how many total handshakes happen? Please explain to me; I`m studying for the Hunter exam that`s tomorrow!!!

OpenStudy (anonymous):

36

OpenStudy (amberlykhan):

How`d you do that?

OpenStudy (anonymous):

6 people shake hands with 6 people 6 x 6 is 36

OpenStudy (amberlykhan):

Thanks!

OpenStudy (anonymous):

no prob

OpenStudy (ranga):

Two people are needed to shake hands. In how many ways can two people be chosen out of 6? 6C2 ways. 6C2 = 6! / {4!2!} = 6.5/2 = 15. Another way to do this problem: Assume the 6 people are standing in a row. The first person shakes hands with 5 people down the line. The second person shakes hands with 4 people down the line, etc. So 5+4+3+2+1 = 15.

OpenStudy (anonymous):

hey its actually 30

OpenStudy (anonymous):

sorry